OpenStudy (anonymous):

Which polynomial is a perfect square trinomial? 49x2 − 28x + 16 4a2 − 20a + 25 25b2 − 20b − 16 16x2 − 24x − 9

OpenStudy (helder_edwin):

you have to complete (if possible) the squares! or take the square roots of the extremes multiply them and then double what you get, for the expression to be a perfect square what u get should be equal to the middle term

OpenStudy (helder_edwin):

for instance in the first option \[ \sqrt{49x^2}=7x \] \[ \sqrt{16}=4 \] then \[ 2(4)(7x)=56x \]

OpenStudy (anonymous):

lol - thats what helper said!

OpenStudy (helder_edwin):

plus when you expand this \[ (a-b)^2=a^2-2ab+b^2 \] the last term if ALWAYS positive. so which ones would you disscard inmediately?